Convert each part of the ratio to percentage:  3 : 1 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have to find total parts by adding the given ratio = 3 + 1 = 4
1stpart = 3⁄4
= (3⁄4) × 100 %
= 3 × 25%
= 75%
2ndpart = 1⁄4
= (1⁄4) × 100%
= 1 × 25
= 25%

Thanks